It is found that the resistance of a coil of wire increases from 40 ohm at 15°C to 50 ohm at 60°C. Calculate the resistance temperature coefficient at 0°C of the conductor material.

1
Expert's answer
2021-12-09T09:23:05-0500

"R_1 = 40 \; ohm \\

R_2 = 50 \; ohm \\

T_1 = 15 \;°C \\

T_2 = 60 \; °C \\

R = R_0(1+αΔT) \\"

R0 = resistance at 0 °C

α = resistance temperature coefficient

"R_1 = R_0(1+α(T_1-0)) \\

40 = R_0(1+15 α) \\

50 = R_0(1+60 α) \\

\frac{40}{50} = \frac{1+15α }{1+60α } \\

4 + 240 α = 5 + 75 α \\

165 α = 1 \\

α = 6.06 \times 10^{-3} \; 1/°C"